OpenStudy (anonymous):

54/9 divided by (41/5 + 81/6) =

OpenStudy (anonymous):

54/9 diveded by 10/1 = 1.66/1 i think

OpenStudy (anonymous):

54/9 = 6 41/5+81/6 = ((41*6)+(81*5))/6*5 = 217/10 (6/1)/(217/10) = 6*10/217*1 = 60/217